Three capacitors, C1-5.00 AuF, C2-2.00 AμF, and C3 7.00 AuF, are connected together. Find the effective capacitance of the group for the following situations. (a) if they are all in parallel AuF (b) if they are all in series
Answer #1

Anonymousanswered this

in parallel

C= C1+C2+C3

=14 μF

in series

C = C1C2C3 / C2C3 +C3C1+C1C2

= 70 / 14 +35 +10

= 70 / 59

=1.186 μF